OpenStudy (anonymous):

Sketch the right triangle and find the length of the side not given. If necessary, approximate the length to the nearest thousandth. leg = 15, hypotenuse = 25

OpenStudy (anonymous):

x^2+15^2= 25^2 x=20

OpenStudy (mertsj):



OpenStudy (mertsj):

Solution just like the Sergeant said.